# Overcoming Reward Model Noise in Instruction-Guided Reinforcement Learning

### v1

@@ Abstract
Vision-language models (VLMs) have gained traction as auxiliary reward models to provide more informative reward signals in sparse reward environments. However, our work reveals a critical vulnerability of this method: a small amount of noise in the reward signal can severely degrade agent performance. In challenging environments with sparse rewards, we show that reinforcement learning agents using VLM-based reward models without proper noise handling perform worse than agents relying solely on exploration-driven methods. We hypothesize that false positive rewards – where the reward model incorrectly assigns rewards to trajectories that do not fulfill the given instruction – are more detrimental to learning than false negatives. Our analysis confirms this hypothesis, revealing that the widely used cosine similarity metric, when applied to comparing agent trajectories and language instructions, is prone to generating false positive reward signals. To address this, we introduce BiMI (Binary Mutual Information), a novel noise-resilient reward function. Our experiments demonstrate that, BiMI significantly boosts the agent performance, with an average improvement ratio of 44.5% across diverse environments with learned, non-oracle VLMs, thereby making VLM-based reward models practical for real-world applications [^1] Code available at https://anonymous.4open.science/r/Official-Overcoming-Reward-Noise-in-Instruction-Guided-RL-E9B6.
@@ 1 Introduction
In real-world reinforcement learning (RL) applications, reward signals are often absent, sparse, or difficult to define explicitly (Clark 2016). To address this, natural language instructions have been adopted as an effective source of reward signals, providing a direct, goal-oriented guidance to agents. This approach, known as instruction-guided reinforcement learning, employs vision-language models (VLMs) to measure the semantic similarity between an agent’s behavior and the provided instructions and reward the agents based on the similarity scores (Kaplan et al. 2017; Goyal et al. 2019, 2020; Du et al. 2023). Specifically, the VLM generates vector embeddings for both the agent’s trajectory history (a sequence of visual frames) and the instructions within a shared embedding space. Notably, prior approaches have consistently adopted cosine similarity to quantify the semantic similarity between vision and language, a metric first demonstrated and proven effective by Frome et al. (2013).
@@
Figure 1: Illustration of false positives in a VLM embedding space. The undesired agent’s trajectory (dashed line) may exhibit high cosine similarity to the instruction (solid line) in the embedding space, as indicated by the proximity of their endpoints in the embedding space. Despite this apparent similarity, these trajectories fundamentally fail to fulfill the intended instruction. Therefore, the rewards guides the agent towards incorrect behaviors. The figure shows three distinct cases of false positives. Please refer to Section 4 for more details.
However, a crucial challenge in VLM-based reward models has been largely overlooked: these models are learned from data rather than being explicitly defined. Consequently, they inherently introduce reward noise due to imperfections in the learning process. This noise manifests in two forms: false positive rewards, where the model incorrectly assigns high rewards to trajectories that do not fulfill the given instructions, and false negative rewards, where the model assigns inadequate rewards to the matched pairs.
Previous research has largely celebrated the potential of VLM reward models under controlled settings, neglecting the pervasive issue of noisy signals. Also, investigations into policy learning failures with learned reward models have typically attributed these failures to false negative rewards and poor training data quality (Du et al. 2023). Our study, however, shifts the focus to the equally critical but less discussed issue of false positive rewards. We argue that (1) false positive rewards are more prevalent and detrimental to learning than false negatives; and (2) using cosine similarity metric is inadequate for capturing the true semantic similarity between trajectories and instructions in RL’s sequential decision-making context, resulting in false positive rewards. As illustrated in Figure 1, this inadequacy stems from two fundamental issues: state entanglement, where the metric fails to account for the impact of actions on state transitions, often assigning high similarity based solely on lexical similarities; and composition insensitivity, where the metric fails to adequately penalize incorrect execution sequences and incomplete task execution.
Recognizing that false positive rewards in VLM-based reward models can significantly hinder performance, we introduce BiMI, a novel reward function designed to mitigate their impact. Specifically, BiMI leverages binary reward signals to directly reduce the occurrence of false positives and incorporates a mutual information term to prevent overfitting to potentially noisy signal sources. Our main contributions are as follows: (1) We provide a detailed analysis of how false positive rewards are generated and its impact on RL, challenging the prevailing assumption that the degeneracy of instruction-guided RL is due to false negative rewards. (2) We introduce BiMI, a noise-robust reward function that significantly improves instruction-guided RL, achieving an average 44.5% performance increase across diverse environments.
@@ 2 Related Work
Use of VLMs as Reward Models. VLMs have been pivotal in robotics studies by serving as reward models that guide agents to follow instructions (Wang et al. 2018; Shridhar et al. 2022; Mahmoudieh et al. 2022). These studies primarily focus on leveraging VLMs t overcome the challenge of manual reward design for complex tasks, without explicitly considering the impact of reward noise or its implications for policy convergence rates. VLM-based reward models have also been commonly used to solve sparse reward environments such as Montezuma’s Revenge in Atari games (Kaplan et al. 2017; Goyal et al. 2019). However, their approaches, which are tested on simplified sub-tasks or require extensive prior knowledge of the environment, inadvertently evade the impact of noisy reward signal. Simplified sub-tasks prevent the accumulation of noise over longer horizons, while prior knowledge about the environment allows agents to reduce their dependence on the noisy reward signal.
Some work sidesteps the noisy reward problem by constructing oracle VLMs, which use internal state information from the game engine rather than pixel information to calculate semantic similarity (Du et al. 2023; Wang et al. 2023). Despite these efforts, the issue of noisy rewards remains a significant challenge in VLM-based reward models. For example, Du et al. (2023) reported difficulties in policy learning with a vanilla learned VLM, attributing the failure to false negatives – instances where the VLM fails to recognize matched pairs. They claimed that an average false negative rate of 11% led to unsuccessful training. However, our research challenges this emphasis on false negatives as the primary obstacle in instruction-guided RL. Our findings indicate that false positive rewards are actually more prevalent and detrimental to the learning process. We provide a comprehensive analysis of noise in VLM-based reward models and examine the comparative impact of false positives and false negatives on the learning efficacy of agents.
Reward Signal from Human Preference. Recent work on Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022) also leverage human preference as a reward signal. However, our work differs in key aspects. Unlike RLHF’s focus on textual outputs, our approach involves evaluating cross-modal similarities between visual and textual data in environments that require long-horizon decision-making and frequent embodied interactions, a domain not typically covered by RLHF.
@@
Mitigating Reward Model Failures. Research by Ghosal et al. (2022) and Fu et al. (2024) has introduced methods to counteract unreliable reward signals from learned VLMs. These strategies involve employing a parallel exploration-based policy alongside the reward-maximizing policy, thereby reducing reliance on potentially misspecified VLM rewards. Our work extends these efforts by providing a comprehensive analysis of noise in VLM-based reward models, with a particular focus on the distinct impacts of false positive and false negative rewards. Furthermore, instead of relying on exploration policies to escape local optima induced by noisy VLM rewards, we directly address the noise within the reward function itself, preventing these pitfalls from arising in the first place.
3 Background
@@
In the context of using instructions as auxiliary guide for RL agents, we frame our task as a MDP defined by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{T},s_{0},r^{\mathit{e}},\gamma\rangle$, where $\mathcal{S}$ represents a set of states $s\in\mathcal{S}$, $\mathcal{A}$ represents a set of actions $a\in\mathcal{A}$, and $\mathcal{T}(s^{\prime}\mid s,a)$ describes the dynamics of the environment. $s_{0}\in\mathcal{S}$ is the initial state, and $\gamma\in[0,1]$ is the reward discount factor. $r^{\mathit{e}}(s,a,s^{\prime})$ is the environmental reward function. In this work, we focus on a sparse reward setting, where the agent receives a non-zero reward only when reaching goal states $S_{G}\subset\mathcal{S}$, and 0 otherwise, with $|S_{G}|\ll|\mathcal{S}|$. An agent’s trajectory is a sequence of states and actions $\tau=\langle s_{t},a_{t},\ldots,s_{t+T}\rangle$, where $t$ is the initial time step and $T$ is the length of the trajectory. The objective is to learn a policy $\pi$ that maximizes the expected cumulative reward of the trajectories $\tau^{\pi}=\{\tau_{1},...,\tau_{n}\mid\tau_{i}\text{ induced by }\pi\}$.
A sparse reward function only realizes a set of acceptable policies, rather than distinguishing between them more finely (as derived from Abel et al. (2021); see Appendix A.1 for details). Building on this, we define the following:
(1) Policy Universe, $\Pi$: the set of all possible policies for the MDP.
(2) Acceptable Policies, $\Pi_{G}$: the policies in $\Pi_{G}$ are those policies whose start-state values $V^{\pi}(s_{0})$ are within a small $\epsilon$ of the optimal value, while all other policies have strictly lower values.
(3) Unacceptable Policies, $\Pi_{B}$: formally, $\Pi_{B}=\Pi\setminus\Pi_{G}$, which refers to the set of policies that do not meet the near-optimal criteria and are therefore considered not acceptable.
A task instruction $L$ is a sequence of natural language sentences $L=\{l_{1},l_{2},\ldots,l_{n}\}$, where each $l_{k}$ induces a set of desired sub-trajectories $\tau^{l_{k}}$. A complete task execution is represented by a trajectory that concatenates instructions’ corresponding sub-trajectories in order, traversing from the initial state $s_{0}$ to a goal state in $S_{G}$.
An auxiliary VLM-based reward model provides more frequent rewards by evaluating the semantic similarity between the agent’s trajectory and the specified instruction sentence. This can be formally represented as $r^{v}(\tau,l_{k})=p(l_{k}|\tau)$ (see Figure 2), where $r^{v}$ denotes the VLM-based reward function. The use of non-Markovian reward functions in MDP environments has been well-established in the literature, particularly through the work on reward machines (Icarte et al. 2018; Corazza et al. 2022). Such a VLM-based reward function consists of two components: an embedding encoder $\phi(\cdot)$ and a similarity function. The reward is calculated as the cosine similarity between the embeddings, which can be expressed as $\frac{\phi(\tau)\cdot\phi(l_{k})}{\|\phi(\tau)\|\|\phi(l_{k})\|}$.
Figure 2: Illustration of reward signals from a vision-language reward model
Reduction of Convergence Rate. To understand the effects of reward noise on policy learning, it’s crucial to analyze its impact on the convergence rate. In sparse reward settings, the convergence rate – defined as the number of training iterations needed to learn an $\epsilon$-optimal policy – can be significantly affected by false positive rewards that are incorrectly given for policies that do not lead to the goal states.
To illustrate this, we analyze the convergence rate by measuring the probability of learning acceptable policies ($P(\pi\in\Pi_{G})$) over each update iteration. The following theorem demonstrates how false positive rewards impact the convergence rate:
Theorem 3.1 (Convergence rate reduction).
In Actor-Critic algorithm, gradient ascent on $Q(s,a)\pi_{\theta_{i}}(a|s)$ pushes the next updated policy $\pi_{\theta_{i+1}}$ in the direction provided by the $Q$ value function. In the direction of maximizing $P(\pi\in\Pi_{G})$, the gradient of $Q(s,a)\pi(a|s)$ can be expressed as follows:
$\nabla_{\theta}Q_{\phi}(s,a)\pi_{\theta}(a|s)=\textit{const}\cdot\mathbb{E}[G^{\pi\in\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$\qquad+(\mathbb{E}[G^{\pi\in\Pi_{G}}]-\mathbb{E}[G^{\pi\in\Pi_{B}}])P(\pi\in\Pi_{G})\nabla_{\theta}\pi_{\theta}(a|s)$
where $\Pi_{G}$ is the set of acceptable policies; $\Pi_{B}$ is the set of unacceptable policies that follow instructions but fail to reach the goal; $\mathbb{E}[G^{\pi}]$ is expectation of the cumulative rewards by executing policy $\pi$ in one episode.
The proof is in Appendix A.2. In this context, false positive rewards correspond to rewards assigned to policies within the set $\Pi_{B}$, which represents undesirable policies. Therefore, $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ can be interpreted as the expected cumulative false positive rewards. This theorem highlights how the presence of $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ reduce the convergence rate by: (1) introducing deviation in the policy update directions (the first term), and (2) reducing the gradient of the desired policy update (the second term). Both effects can significantly hinder the convergence process.
The issue of noisy rewards is not merely a theoretical concern; it has practical implications for the effectiveness of instruction-guided RL in real-world applications. In the next section, we will investigate how cosine similarity metric used in VLM-based reward models inevitably introduce false positive rewards. Additionally, we will empirically evaluate the impact of both false positives and false negatives on the learning efficacy of agents in sparse reward environments.
4 Approximation Error of Cosine Similarity

### v4

# The Dark Side of Rich Rewards: Understanding and Mitigating Noise in VLM Rewards

@@ Abstract
While Vision-Language Models (VLMs) are increasingly used to generate reward signals for training embodied agents to follow instructions, our research reveals that agents guided by VLM rewards often underperform compared to those employing only intrinsic (exploration-driven) rewards, contradicting expectations set by recent work. We hypothesize that false positive rewards – instances where unintended trajectories are incorrectly rewarded – are more detrimental than false negatives. We confirmed this hypothesis, revealing that the widely used cosine similarity metric is prone to false positive estimates. To address this, we introduce BiMI (Binary Mutual Information), a novel reward function designed to mitigate noise. BiMI significantly enhances learning efficiency across diverse and challenging embodied navigation environments. Our findings offer a nuanced understanding of how different types of reward noise impact agent learning and highlight the importance of addressing multimodal reward signal noise when training embodied agents [^1] Code available at https://shorturl.at/VsH70.
The Dark Side of Rich Rewards:
Understanding and Mitigating Noise in VLM Rewards
Sukai Huang1, Shu-Wei Liu2, Nir Lipovetzky1 and Trevor Cohn1thanks: Now at Google DeepMind 1School of Computing and Information Systems, The University of Melbourne, Australia 2 Max Planck Institute for the Physics of Complex Systems, Nöthnitzer Strasse 38, 01187 Dresden, Germany sukaih@student.unimelb.edu.au, sliu@pks.mpg.de, {nir.lipovetzky, trevor.cohn}@unimelb.edu.au
@@ 1 Introduction
Natural language instructions are increasingly recognized as a valuable source of reward signals for guiding embodied agents to learn complex tasks. In particular, a growing trend in embodied agent learning involves using vision-language models (VLMs) for reward modeling. This approach measures the semantic similarity – often quantified by cosine similarity – between the embedding representations of an agent’s behaviors (i.e., past trajectories) and the provided instructions, all within the same embedding space (Kaplan et al., 2017; Goyal et al., 2019, 2020; Du et al., 2023).
@@
Figure 1: Schematic diagram of false positives in a VLM embedding space. The unintended agent’s trajectory (dashed line) may exhibit high cosine similarity to the instruction (solid line) in the embedding space, as indicated by the proximity of their endpoints in the embedding space. Despite this apparent similarity, the unintended trajectory fails to fulfill the instruction. This illustrates how approximation errors can lead to false positive rewards. Refer to Section 5 for more details.
However, we observed that embodied reinforcement learning (RL) agents trained with VLM rewards, while effective in simplified settings, often struggled with tasks involving complex dynamics and longer action horizons. This is evident in several recent works – for instance, Goyal et al. (2019) reported the effective use of VLM rewards in Montezuma’s Revenge, a notoriously challenging Atari game. However, we observed that this success was confined to individual sub-tasks and the agent struggled when attempting to scale up to the full game. Similarly, Du et al. (2023) demonstrated impressive performance of VLM rewards in guiding agents within a 2D survival game. However, their study was conducted in a modified environment with a reduced observation and action space using internal game state information and manually defined macro actions. Consequently, when tested in the original, unmodified environment, their agent’s performance did not exceed that of agents using only intrinsic (exploration-driven) rewards.
The consistent underperformance of VLM rewards, particularly their unexpected failure to outperform intrinsic rewards, raised concerns about their reliability. This discrepancy, where VLM rewards underperformed contrary to their perceived potential, prompted us to investigate the underlying causes of this performance gap. Our findings indicate that noisy reward estimates in VLMs are a key factor contributing to poor learning efficacy. Specifically, our analysis centered around two classes of noise: false positives, which involve rewarding unintended trajectories, and false negatives, which occur when correct trajectories are not rewarded. We posit that false positive rewards (see Figure 1) are not only more prevalent but potentially more detrimental to the learning process than false negatives, a hypothesis supported by our empirical and theoretical findings. Among various sources of reward noise, our study particularly investigates the approximation errors within the commonly used cosine similarity metric. We examine how these errors generate false positive rewards, which in turn hinder learning.
To this end, we propose a novel reward function, BiMI (Binary Mutual Information Reward). It uses binary reward signals to directly reduce the occurrence of false positives and incorporates a mutual information term to prevent overfitting to noisy signal sources. Our experiments demonstrate that BiMI significantly improves the learning efficacy of instruction following agents trained by VLM rewards across various challenging environments.
@@ 2 Related Work
Using VLMs as Reward Models. VLMs have been pivotal in robotics studies, serving as reward models that guide agents to follow instructions (Wang et al., 2018; Shridhar et al., 2022; Mahmoudieh et al., 2022). While most research primarily focuses on leveraging VLMs to overcome the challenge of manual reward design for complex tasks (Clark, 2016), the impact of reward noise and its implications for policy convergence rates are often overlooked. As mentioned in Section 1, some work sidesteps the noisy reward problem by accessing internal state information from the game engine as well as providing predefined action macros (Du et al., 2023; Wang et al., 2023), thereby preventing the accumulation of noise over longer horizon. However, understanding the impact of reward noise from VLMs is crucial for developing a reliable language interface for embodied learning agents in real-world applications.
@@
Mitigating Reward Model Failures. Research by Ghosal et al. (2022) and Fu et al. (2024) has introduced methods to counteract unreliable rewards from learned VLMs. These strategies involve employing a parallel exploration-based policy alongside the reward-maximizing policy, thereby reducing reliance on potentially misspecified VLM rewards. Our work contributes to this growing body of research by proposing a novel reward function that directly mitigates the impact of false positive rewards from VLM-based models, complementing approaches that use exploration policies to escape local optima. Furthermore, we tested the synergy of combining our reward function with exploration strategies, demonstrating how these approaches can be integrated for further advancements.
3 Formal Problem Statement
@@
We frame our task as an MDP defined by a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},r^{e},\gamma\rangle$, where $\mathcal{S}$ represents a set of states $s\in\mathcal{S}$, $\mathcal{A}$ represents a set of actions $a\in\mathcal{A}$, and $\mathcal{P}(s^{\prime}|s,a)$ describes the dynamics of the environment. $s_{0}\in\mathcal{S}$ is the initial state and $\gamma\in(0,1)$ is the reward discount factor. $r^{e}(s,a)$ is the environmental reward function. An agent’s trajectory is a sequence of states and actions $\tau_{t}=\langle s_{0},a_{0},\ldots,s_{t}\rangle$.
In this work, we focus on a sparse reward setting, where the agent receives a $+1$ reward only when reaching goal states $S_{G}\subset\mathcal{S}$, and 0 otherwise, with $|S_{G}|\ll|\mathcal{S}|$. This sparse reward setting motivates the use of expert instructions and VLMs to provide auxiliary reward signals for more effective RL. Specifically, we have a walkthrough $L$ that breaks down a complex task into $n$ expert-defined sub-tasks, each represented by a natural language instruction that is not necessarily atomic and can encompass multiple finer sub-goals ($L=\{l_{1},l_{2},\ldots,l_{n}\}$). By following these sequential instructions, the agent can navigate from the initial state towards the goal states. A dedicated non-Markovian VLM-based reward model $r^{v}(\tau_{t},l_{m(t)})$ is used to assess how well the agent’s trajectory at current time $t$ fulfills an instruction sentence $l_{m(t)}$. Here, $m(t)$ is a pointer at time step $t$ that indicates the current instruction the agent is trying to complete [^2] The use of non-Markovian reward functions in MDP has been well-established, particularly through the work on reward machines (Icarte et al., 2018; Corazza et al., 2022). For a complete evaluation, we also tested Markovian version of VLM reward function (Pixl2R) in our experiments.. The VLM provides auxiliary rewards by evaluating the semantic similarity between $\tau_{t}$ and $l_{m(t)}$, as illustrated in Figure 2 (see Appendix A.1 for details).
Figure 2: Illustration of reward from a VLM
4 Theoretical Analysis
In this section, we theoretically show how auxiliary rewards influence policy convergence. Initially, we establish that providing auxiliary rewards that reflect progress towards the goal state typically accelerates the convergence process compared to scenarios where only sparse environmental reward $r^{e}$ is used. Then, we show that the presence of false positive rewards has a more detrimental effect on convergence than false negatives.
4.1 Auxiliary Rewards and Convergence
With sparse rewards, the gradient landscape is nearly flat, making gradient-based updates indistinguishable from a random walk in parameter space (Antognini and Sohl-Dickstein, 2018). In this problem, we are interested in $D:=\|\theta_{goal}-\theta_{0}\|$ which is the distance in parameter space from initial parameters $\theta_{0}$ to goal parameters $\theta_{goal}$. We make the following assumption
Assumption 4.1.
Expert knowledge can guide the parameter search along a path in parameter space, defined by a sequence of $n$ intermediate parameter vectors $\theta_{1},\dots,\theta_{n}$, where each $\theta_{i}$ represents the parameters after learning sub-task $l_{i}$. As a result, the overall distance $D$ can be decomposed into segments: $D\approx\sum_{k=1}^{n-1}d_{i}$, where $d_{i}=\|\theta_{i+1}-\theta_{i}\|$.
We also prove the following proposition
Proposition 4.2.
The sum of expected time for a series of random walks, each covering the shorter distance of an individual sub-task, is less than the expected time to travel the entire distance $D$ in one long random walk:
$\frac{1}{n-1}\mathbb{E}[T_{D}]\leq\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]<\mathbb{E}[T_{D}].$
and thus show that subgoal-based auxiliary rewards improve the convergence of random walk optimization in a sparse reward landscape up to a factor of $(n-1)$ with details in Appendix A.2.1.
4.2 Connection to Heuristic-Guided RL
The problem and the possible solutions can be framed within the context of Heuristic-Guided Reinforcement Learning (HuRL) by Cheng et al. (2021). HuRL mandates that auxiliary reward signals serve as heuristics, where $h:\mathcal{S}\rightarrow\mathbb{R}$ approximates the future total rewards an agent expects to get starting from state $s$ under the optimal policy $\pi^{*}$ (i.e., $h(s)\approx V^{*}(s)$). These heuristics typically come from domain knowledge, aligning with expert instructions. We make the following assumption
Assumption 4.3.
The fulfillment of $l_{m(t)}$ can be captured by intermediate goal states $\mathcal{S}^{m(t)}$.
This assumption is justified because often instructions in VLM come with end goals that can be indicated by definite states. For further discussions, please refer to Section 9. This assumption is required so that we can construct a domain-knowledge Markovian heuristic reward based on non-Markovian VLM rewards (see Appendix A.1.3 for the exact formula). With this construction, we are estimating the possible reward in the future by evaluating how well the agent has followed expert instruction in the past. It is consistent with the intuition of a heuristic reward based on semantic similarity because the better the agent has been following expert instructions, the more likely it is on the right track to collect more future rewards.
Using HuRL framework allows us to analyze how false positive rewards influence on policy performance gap, defined as $V^{*}(s_{0})-V^{\pi}(s_{0})$. We will demonstrate that false positive rewards increase the upper bound of this gap, whereas false negative rewards maintain this upper bound.
4.3 False Positives vs. False Negatives
To begin, we provide the formal definition of false positive auxiliary rewards from both instruction-following (IF) and heuristic perspectives:
Definition 4.4 (False Positive Rewards).
A false positive reward occurs when:
IF Perspective: For a trajectory $\tau_{t}$ that does not satisfy instruction $l_{m(t)}$, the VLM-based reward $r^{v}(\tau_{t},l_{m(t)})$ ranges between 00 and $1$.
Heuristic Perspective: The heuristic $h(s_{t})>V^{*}(s_{t})$, overestimating the optimal value of $s_{t}$.
Proposition 4.5.
False positive rewards in the IF perspective imply false positive rewards in the heuristic perspective.
The heuristic is approximated as $h(s_{t})\approx V^{*}(s_{t})=1\cdot\gamma^{\widetilde{T}-t}$, calculated based on an assumed optimal path length $\widetilde{T}$. Here, $\widetilde{T}-t$ measures the remaining steps towards the goal. When a trajectory $\tau_{t}$ receives a high reward but fails to fulfill instruction $l_{m(t)}$, it corresponds to a high $h(s_{t})$, thus a low $\widetilde{T}-t$. However, since the agent must eventually redo $l_{m(t)}$, $s_{t}$ is actually further from the goal than estimated. Therefore, the actual distance $T-t$ to reach the goal will exceed $\widetilde{T}-t$, and $V^{*}(s_{t})$, calculated with the actual $T$, will be smaller than $h(s_{t})$, derived from $\widetilde{T}$. This thus explains how a false positive reward from an instruction-following (or VLM) perspective leads to an overestimation of the heuristic. Researchers have advocated the benefits of pessimistic value estimation to enhance the stability of RL algorithms (Kumar et al., 2020; Jin et al., 2021). In HuRL, Cheng et al. (2021) further identify a beneficial property: when a heuristic is pessimistic with respect to $\mathcal{M}$, it results in a smaller upper bound on the performance gap.
Definition 4.6 (Pessimistic $h$).
Let Bellman operator $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$. A heuristic function $h$ is said to be pessimistic with respect to an MDP $\mathcal{M}$ if $\max_{a}(\mathcal{B}h)(s,a)\geq h(s)$. This condition essentially means that the heuristic $h$ never overestimates the true value of a state.
Proposition 4.7.
Even if the heuristic remains conservative for all successor states, a single overestimation ($h(s)>V^{*}(s)$) can violate the pessimistic condition by causing $max_{a}(\mathcal{B}h)(s,a)<h(s)$.
Informally, our result on the impact of false positive/negative rewards on convergence is as follows:
Theorem 4.8.
Borrowing from Cheng et al. (2021), the performance gap in RL when using heuristics can be broken down into a regret term and a bias term, where false negatives maintain the upper bound of bias while false positives increase the bias by breaking the heuristic’s pessimism, thereby potentially leading to slower convergence.
See Appendix A.2.2 for formal version and detailed proofs. In the next section, we present a case study on cosine similarity metrics, demonstrating how they contribute to false positive rewards in learned reward models.
5 False Positives From Cosine Similarity